Required information (The following information applies to the questions displayed below.] Sweeten Company had no jobs in pro2. How much manufacturing overhead was applied from the Molding Department to Job P and how much was applied to Job Q? (Do no3. How much manufacturing overhead was applied from the Fabrication Department to Job P and how much was applied to Job Q? (D4. What was the total manufacturing cost assigned to Job P? (Do not round intermediate calculations.) Total manufacturing cos5. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answ6. What was the total manufacturing cost assigned to Job Q? (Do not round intermediate calculations.) Total manufacturing cos7. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final answ8. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establi9. What was Sweeten Companys cost of goods sold for March? (Do not round intermediate calculations.) Cost of goods sold10. What was the companys plantwide predetermined overhead rate? (Round your answer to 2 decimal places.) Predetermined over11. How much manufacturing overhead was applied to Job P and how much was applied to Job Q? (Do not round intermediate calcul12. If Job P included 20 units, what was its unit product cost? (Do not round intermediate calculations. Round your final ans13. If Job Q included 30 units, what was its unit product cost? (Do not round intermediate calculations. Round your final ans14. Assume that Sweeten Company used cost-plus pricing (and a markup percentage of 80% of total manufacturing cost) to establ15. What was Sweeten Companys cost of goods sold for March? (Do not round intermediate calculations.) Cost of goods sold
Answer #1

1) Molding overhead rate = (13750/2500+2.9) = 8.4 per MH

Fabrication overhead rate = (17250/1500+3.7) = 15.20 per MH

2) Molding overhead applied

Job P Job Q
Molding overhead applied 3200*8.4 = 26880 2300*8.4 = 19320

3) Fabrication overhead applied

Job P Job Q
Fabrication overhead applied 2100*15.20 = 31920 36480

4) Total manufacturing cost for Job P

Direct material 28000
Direct labor 33000
Molding overhead 26880
Fabrication overhead 31920
Total manufacturing cost 119800

5) Unit product cost = 119800/20 = 5990

Note: Pls post whole que in the group of 4-4 sub parts as per HOMEWORKLIB RULES